A monopoly that faces a demand curve given by q= 1-P and has a constant marginal cost is 0.1 In this situation the monopoly’s profit maximizing output level is
1
Expert's answer
2021-10-25T08:48:14-0400

P = 1 - Q

MR = 1 - 2Q

MC = 0.1

MR = MC

1 - 2Q = 0.1

1 - 0.1 = 2Q

0.9 = 2Q

Q = 0.45